OpenStudy (anonymous):

Solve the system by the method of substitution. 2x^2 + 4x − y = 19 2x − y = −5

hero (hero):

Hang on a sec

OpenStudy (anonymous):

okay

hero (hero):

The goal here is to get one of the equations all in terms of one variable. If you managed to get the first equation in terms of x only, it would simplify to a quadratic equation. From there you using factoring methods.

OpenStudy (anonymous):

im suppose too find two (x,y) but I got it wrong.

hero (hero):

Re-writing the second equation, y = 2x + 5 You can substitute the expression for y into the first equation to get: 2x^2 + 4x - (2x + 5) = 19

hero (hero):

We will find the (x,y) pair I assure you. But first, we needed to replace y with an appropriate expression for x that would work.

OpenStudy (anonymous):

i got 2x^2+2x-24=0

OpenStudy (anonymous):

then factor for x?

hero (hero):

Well, if you factor out 2 on the left side first, you'll get 2(x^2 - x - 12) = 0 Then divide both sides by 2 to get x^2 - x - 12 = 0 That should make it easier to factor.

OpenStudy (anonymous):

x=-3 and4?

hero (hero):

x^2 - x - 12 = 0 x^2 -4x + 3x -4*3 = 0 x(x - 4) + 3(x - 4) = 0 (x - 4)(x + 3) = 0

OpenStudy (anonymous):

so now i plug in 4 and -3 to find y?

OpenStudy (anonymous):

i got (4,13) and (-3,-1) and it was wrong

hero (hero):

I made a sign mistake x^2 + x - 12 = 0 Factor that

OpenStudy (anonymous):

\[(x-3) (x+4)=0 \]